# Massive Fields Affected by Echoes: New Physics vs. Astrophysical Environment

### v1

@@ I Introduction
Echoes are modifications of a signal at late times stipulated by the secondary scatterings of the wave from additional maxima of the effective potential Cardoso et al. (2016a); Cardoso and Pani (2017); Barausse et al. (2014). While echoes of massless fields have been studied in a great number publications (see, for instance, Cardoso et al. (2016a); Cardoso and Pani (2017); Barausse et al. (2014); Huang et al. (2022); Konoplya et al. (2019); Abedi et al. (2017); Mark et al. (2017); Wang et al. (2020); Bronnikov and Konoplya (2020); Bueno et al. (2018); Churilova et al. (2021); Guo et al. (2022); Churilova and Stuchlik (2020); Li and Piao (2019); Buoninfante (2020); Liu et al. (2021); Chowdhury et al. (2022); Yang et al. (2024a); Shen et al. (2024); Yang et al. (2024b) and references therein), echoes in the massive gravity, to the best of our knowledge, were discussed only in Dong and Stojkovic (2021) where no effects for effectively massive fields were reported. However, the time-domain profiles for the Kaluza-Klein gravitational degrees of freedom in the braneworld scenario have been recently obtained in Tan et al. (2024).
@@
At the same time, there are several motivations to study the evolution of perturbations in massive fields. Firstly, fields that are initially massless may acquire an effective mass term due to the influence of tidal forces in brane-world models and other extra-dimensional scenarios Seahra et al. (2005); Ishihara et al. (2008), or in the presence of a magnetic field Konoplya and Fontana (2008); Konoplya (2008); Wu and Xu (2015). Additionally, massive gravitons and other massive particles are expected to contribute to long-wavelength gravitational waves Konoplya and Zhidenko (2024a), which are currently being investigated through Pulsar Timing Array observations Afzal et al. (2023).
@@
The quasinormal modes of massive fields differ qualitatively from those of massless fields, particularly due to the strong suppression of the damping rate. In some cases, this leads to the existence of arbitrarily long-lived modes, known as quasi-resonances, within the spectrum Ohashi and Sakagami (2004); Konoplya and Zhidenko (2005, 2006); Zhidenko (2006). Moreover, the asymptotic tails of massive fields exhibit distinct behavior, as they do not decay according to the power-law characteristic of massless fields. Instead, they oscillate and decay slowly with a power-law envelope (see the review in Konoplya and Zhidenko (2011)).
@@
In this work, we address a gap in the literature by studying echoes in massive fields. To this end, we consider three different models: the simplest massive scalar field in a Schwarzschild background, the squashed Kaluza-Klein black hole, and a wormhole mimicking the Schwarzschild solution Damour and Solodukhin (2007). In the first two models, echoes are induced by a deformation in the effective potential, creating a bump. In contrast, the third model features a double-well potential, producing echoes without the need for an additional bump.
@@
We have demonstrated that echoes in massive fields are markedly different from those in massless fields. Specifically, when the bump is localised in the far zone, they affect the signal more during the asymptotic tail stage, leading to the oscillation of the power law envelope. The near-horizon bumps, one the contrary, induce the distinctive echoes effects in some cases at the earlier stage of quasinormal ringing. Given the slow decay of the asymptotic tails in massive fields, this enhances the significance of echoes in such fields.
@@
The paper is organized as follows. In Section II, we examine the echoes of a massive scalar field in a Schwarzschild background, focusing on cases where the effective potential has a bump both at a distance and near the horizon. In Section III, we analyze the echoes of gravitational perturbations (effectively massive) around squashed Kaluza-Klein black holes. Section IV is dedicated to the study of echoes of massive fields around the Schwarzschild-like wormhole.
@@
II Schwarzschild black hole with the Gaussian bump in the effective potential
@@
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\,,$ (1)
where
$f(r)\equiv 1-\frac{r_{0}}{r},$
and $r_{0}=2GM$ is the radius of the event horizon.
The Klein-Gordon equation for a massive scalar field
$\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\Phi\right)=\mu^{2}\Phi,$ (2)
after separation of the angular part
$\Phi=\frac{1}{r}\Psi(t,r)Y_{\ell}^{m}(\theta,\phi),$
can be reduced to the wavelike form
$\left(\frac{\partial^{2}}{\partial t^{2}}-\frac{\partial^{2}}{\partial r_{*}^{2}}+V(r)\right)\Psi(t,r)=0\,.$ (3)
Here $r_{*}$ is the tortoise coordinate,
$dr_{*}=\frac{dr}{f(r)}\,,$
and the effective potential is given by
$V(r)=f(r)\left(\mu^{2}+\frac{\ell(\ell+1)}{r^{2}}+\frac{f^{\prime}(r)}{r}\right)\,.$ (4)
We add a bump to the effective potential
$V(r)\rightarrow V(r)+\delta V(r),$ (5)
where
$\delta V(r)=Af(r)e^{-(r-r_{m})^{2}/\kappa}.$ (6)
To study the time-domain evolution of perturbations in these black hole models, we use the well-known time-domain integration scheme proposed in Gundlach et al. (1994). We apply the discretization scheme in terms of the light-cone variables $u=t-r_{*}$ and $v=t+r_{*}$,
$\Psi\left(N\right)$ $=$ $\Psi\left(W\right)+\Psi\left(E\right)-\Psi\left(S\right)$
$-\Delta^{2}V\left(S\right)\frac{\Psi\left(W\right)+\Psi\left(E\right)}{8}+{\cal O}\left(\Delta^{4}\right)\,,$
where we used the following notation for the points: $N\equiv\left(u+\Delta,v+\Delta\right)$, $W\equiv\left(u+\Delta,v\right)$, $E\equiv\left(u,v+\Delta\right)$, and $S\equiv\left(u,v\right)$. This discretization scheme has been widely applied in numerous studies (see Konoplya (2024); Dubinsky and Zinhailo (2024); Dubinsky (2024); Konoplya and Zinhailo (2020); Skvortsova (2024); Bolokhov (2024a, b); Malik (2024a, b); Konoplya and Stashko (2024) for recent examples). The Gaussian initial data are imposed on the two null surfaces, $u=0$ and $v=0$:
$\begin{array}[]{rcl}\Psi(u,0)&=&\exp\left(\dfrac{u-p}{2}\right)^{2},\\ \Psi(0,v)&=&\exp\left(\dfrac{v+p}{2}\right)^{2},\\ \end{array}$ (8)
@@
We can observe how the perturbations evolve over time, depending on various parameters such as the bump characteristics, the field’s mass, and the observer’s position. From the plot 1 one can see that, when the bump is located far from the black hole, echoes appear at very late times, modifying the power-law tail envelope, which becomes oscillatory with a non-constant period and amplitude. For smaller bumps, the amplitude of the oscillations decreases, and the farther the bump is from the black hole, the later the echoes appear. When the bump is closer to the event horizon, as shown in Fig. 2, both the intermediate tails and the final stage of the ringing phase are significantly modified. However, the asymptotic late-time tails, being governed by the asymptotic behavior of the potential, change insignificantly. Therefore, in order to observe the distinct echoes for the massive field, we need a significant modification of the effective potential outside the radiation zone. Such a modification can appear at a distance, due to another compact object with a comparable mass, e.g., a black hole or a neutron star, or, alternatively, due to the effects of new physics near the horizon or instead of the horizon. The latter effect for a Schwarzschild wormhole we shall consider in Sec. IV.
@@
Figure 3: Effective potential and time-domain profile for a massive scalar field perturbations around the Schwarzschild black hole with a bump: $r_{0}=1$, $\ell=1$, $r_{m}=8$, $\kappa=1$, $\mu=1$ (blue) and $\mu=0$ (red), $A=1/100$ (left panel) and $A=1/12$ (right panel). The observer is situated at $r_{p}=3$, and the center of the Gaussian wave-package is between the observer and the event horizon, $p=16$.
For sufficiently small bumps near the event horizon, the time-domain profile during the ringdown phase does not differ significantly from that of the spacetime without a bump. This result aligns with Konoplya and Zhidenko (2024b); Konoplya (2023), where it was demonstrated that the fundamental mode typically changes only slightly when the potential is slightly deformed near the horizon, with stronger deviations seen in the first few overtones. These overtones decay quickly, and for massive fields, their contribution to the time-domain profile is suppressed because (a) the fundamental mode decays more slowly and (b) the tails of massive fields begin to dominate earlier, leaving a relatively short period for the ringing phase.
When comparing the echoes for massive and massless fields (see Fig. 3), we observe that the massless field, being long ranged, scatters the energy faster and, therefore the amplitude of echoes are many orders smaller. Even for an unrealistically, for astrophysical environment, large bump ($A=1/12$) we see that, although the amplitude of the first echo is of the same order of magnitude as the intermediate time tail, each next echo’s amplitude becomes by orders smaller than the echo for the massive field.
III Gravitational perturbations of squashed black holes in the Kaluza-Klein theory

### v2

@@ 1 Introduction
Echoes are modifications of a signal at late times stipulated by the secondary scatterings of the wave from additional maxima of the effective potential [1, 2, 3]. While echoes of massless fields have been studied in a great number publications (see, for instance, [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] and references therein), echoes in the massive gravity, to the best of our knowledge, were discussed only in [21] where no effects for effectively massive fields were reported. However, the time-domain profiles for the Kaluza-Klein gravitational degrees of freedom in the braneworld scenario have been recently obtained in [22].
@@
At the same time, there are several motivations to study the evolution of perturbations in massive fields. Firstly, fields that are initially massless may acquire an effective mass term due to the influence of tidal forces in brane-world models and other extra-dimensional scenarios [23, 24], or in the presence of a magnetic field [25, 26, 27]. Additionally, massive gravitons and other massive particles are expected to contribute to long-wavelength gravitational waves [28], which are currently being investigated through Pulsar Timing Array observations [29].
Although the LIGO detection of gravitational-wave signals has placed stringent constraints on the graviton mass $\mu$, or, equivalently, a lower bound on the graviton Compton wavelength $\lambda_{c}$ [30, 31],
$\mu c^{2}\lesssim 2\cdot 10^{-23}eV,\quad\lambda_{c}\gtrsim 1~\mbox{light year},$ (1.1)
these limits apply primarily to theories in which gravity is mediated solely by a single massive spin-2 field. In contrast, a broad class of alternative gravity theories admits both massless and massive gravitational degrees of freedom, with the massive modes typically arising in addition to the standard massless graviton. In such frameworks, the above observational bounds do not necessarily apply, and the massive sector can remain essentially unconstrained (see, e.g., the discussion in [32]).
Massive gravitational degrees of freedom, much like other massive fields, can also affect the evolution of gravitational perturbations through the nonlinear backreaction of the stress–energy carried by the waves, similarly to the gravitational-wave memory in massless gravity [33]. A simplified setting of static perturbations of the massive field has been considered in [21].
@@
The quasinormal modes of massive fields differ qualitatively from those of massless fields, particularly due to the strong suppression of the damping rate. In some cases, this leads to the existence of arbitrarily long-lived modes, known as quasi-resonances, within the spectrum [34, 35, 36, 37, 38, 39, 40, 41]. Moreover, the asymptotic tails of massive fields exhibit distinct behavior, as they do not decay according to the power-law characteristic of massless fields. Instead, they oscillate and decay slowly with a power-law envelope (see the review in [42]).
Detecting these asymptotic tails poses substantial observational challenges: By the time the tail phase begins, signal amplitudes lie many orders of magnitude below the preceding quasinormal ringing and are deeply submerged in detector noise. For massless fields, the decay follows a smooth power-law falloff, quickly rendering the tail undetectable. However, for massive fields, the situation changes significantly: These tails not only decay more slowly than their massless counterparts, but can persist for much longer timescales, extending the observation window in which they might be extracted from noise. The oscillation frequency is determined by the black hole mass: For stellar-mass black holes, the frequency lies within the sensitivity band of ground-based detectors such as Advanced LIGO [43]. For supermassive black holes, the range shifts to the planned bandwidth of space-based antennas such as eLISA [44]. Although their amplitude is typically about $10^{4}$ times smaller than that of the corresponding quasinormal modes, the overlap between the natural frequency of the massive tail and the detector response, combined with the longer-lived nature of the signal, significantly improves the prospects for detection, provided that search algorithms can accommodate the distinctive oscillatory phase structure [45].
@@
In this work, we address a gap in the literature by studying echoes in massive fields. To this end, we consider three different models: the simplest massive scalar field in a Schwarzschild background, the squashed Kaluza-Klein black hole, and a wormhole mimicking the Schwarzschild solution [46]. In the first two models, echoes are induced by a deformation in the effective potential, creating a bump. In contrast, the third model features a double-well potential, producing echoes without the need for an additional bump.
@@
We have demonstrated that echoes in massive fields are markedly different from those in massless fields. Specifically, when the bump is localised in the far zone, they affect the signal more during the asymptotic tail stage, where it induces oscillations in the power-law decay envelope. Such deformations of the effective potential are typically associated with environmental effects, including the presence of accretion disks, nearby companion compact objects, active galactic nuclei, or extended matter distributions such as scalar or dark matter clouds (see, e.g., [5] and references therein). The near-horizon bumps, one the contrary, induce the distinctive echoes effects in some cases at the earlier stage of quasinormal ringing. Given the slow decay of the asymptotic tails in massive fields, this enhances the significance of echoes in such fields.
@@
The paper is organized as follows. In Section 2, we examine the echoes of a massive scalar field in a Schwarzschild background, focusing on cases where the effective potential has a bump both at a distance and near the horizon. In Section 3, we analyze the echoes of gravitational perturbations (effectively massive) around squashed Kaluza-Klein black holes. Section 4 is dedicated to the study of echoes of massive fields around the Schwarzschild-like wormhole.
@@
2 Schwarzschild black hole with the Gaussian bump in the effective potential
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\,$, (2.1)
where
$f(r)\equiv 1-\frac{r_{0}}{r},$
and $r_{0}=2GM$ is the radius of the event horizon.
The Klein-Gordon equation for a massive scalar field
$\frac{1}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\Phi\right)=\mu^{2}\Phi,$ (2.2)
after separation of the angular part
$\Phi=\frac{1}{r}\Psi(t,r)Y_{\ell}^{m}(\theta,\phi),$
can be reduced to the wavelike form
$\left(\frac{\partial^{2}}{\partial t^{2}}-\frac{\partial^{2}}{\partial r_{*}^{2}}+V(r)\right)\Psi(t,r)=0\,.$ (2.3)
Here $r_{*}$ is the tortoise coordinate,
$dr_{*}=\frac{dr}{f(r)}\,$,
and the effective potential is given by
$V(r)=f(r)\left(\mu^{2}+\frac{\ell(\ell+1)}{r^{2}}+\frac{f^{\prime}(r)}{r}\right)\,.$ (2.4)
We add a bump to the effective potential
$V(r)\rightarrow V(r)+\delta V(r),$ (2.5)
where
$\delta V(r)=Af(r)e^{-(r-r_{m})^{2}/\kappa}.$ (2.6)
To study the time-domain evolution of perturbations in these black hole models, we use the well-known time-domain integration scheme proposed in [47]. We apply the discretization scheme in terms of the light-cone variables $u=t-r_{*}$ and $v=t+r_{*}$,
$\Psi\left(N\right)$ $=$ $\Psi\left(W\right)+\Psi\left(E\right)-\Psi\left(S\right)-\Delta^{2}V\left(S\right)\frac{\Psi\left(W\right)+\Psi\left(E\right)}{8}+{\cal O}\left(\Delta^{4}\right)\,$, (2.7)
where we used the following notation for the points: $N\equiv\left(u+\Delta,v+\Delta\right)$, $W\equiv\left(u+\Delta,v\right)$, $E\equiv\left(u,v+\Delta\right)$, and $S\equiv\left(u,v\right)$. This discretization scheme has been widely applied in numerous studies (see [48, 49, 50, 51, 52, 53, 39, 54, 55, 56] for recent examples). The Gaussian initial data are imposed on the two null surfaces, $u=0$ and $v=0$:
[…]
Before discussing the effect of bumps onto the late-time decay of massive fields, let us first review the decay properties without such bumps. The behavior of the late-time field can be divided into two main regimes.
At intermediate late times, after the quasinormal ringing has decayed but before the asymptotic tail dominates, the field decays as a power law:
$\Phi(t,r)\propto t^{-p}\sin(\mu t+\varphi),$ (2.9)
where $\varphi$ is a phase shift and the power index $p$ depends on the multipole number $\ell$ and spacetime properties. In the Schwarzschild background, for an observer at a fixed radius:
$p=\ell+\frac{3}{2}.$ (2.10)
This behavior was first derived analytically in [57, 58, 59] for Schwarzschild and Reissner-Nordström black holes.
At very late times ($\mu t\gg 1/\mu^{2}M^{2}$), the field exhibits an oscillatory inverse power-law tail, commonly known as the asymptotic tail:
$\Phi(t,r)\propto t^{-5/6}\sin(\mu t+\varphi).$ (2.11)
This decay rate is universal, in the sense that it does not depend on the multipole number $\ell$ or spacetime curvature, and it arises from backscattering off the asymptotically flat region. The power 5/6-5/6 was analytically derived in [58] and confirmed numerically in many follow-up works [32]. It is valid for Schwarzschild and other asymptotically flat black holes (such as Reissner-Nordström and slowly rotating Kerr).
Now, keeping in mind the qualitatively different late-time decay of massive fields, we are ready to discuss how potential bumps affect the evolution of perturbations. We can observe how the perturbations evolve over time, depending on various parameters such as the bump characteristics, the field’s mass, and the observer’s position. The modification of the late-time signal (echoes) originates from the partial reflection of the wave by a secondary peak in the effective potential, followed by its subsequent interaction with the black hole’s near-horizon region. Physically, this process corresponds to a portion of the outgoing radiation being trapped between the primary (near-horizon) barrier and the secondary, more distant potential barrier. Each round trip between these barriers generates a delayed, weaker replica of the original signal, giving rise to a sequence of echoes in the waveform. For a massless field, the characteristic delay time marking the onset of the first echo is approximately given by the light travel time from the black hole to the secondary peak and back. Therefore, this delay is approximately equal to twice the separation $\Delta r_{*}$ between the two potential peaks [60],
$\Delta t\simeq 2\Delta r_{*}=2\intop_{3r_{0}/2}^{~r_{m}}\frac{dr}{f(r)}.$ (2.12)
For a massive field, the modification of the late-time behavior also emerges on a similar timescale. In the regime of large field masses, the centrifugal barrier is suppressed, and the potential effectively resembles a step-like structure with a single bump rather than two distinct peaks. In this case, the modification of the late-time tail manifests as amplitude beating. Unlike the regular modulations arising from interference between long-lived quasinormal modes, such as those appearing due to superradiance [61], this beating is irregular. As time progresses, the period of the beating gradually increases, leading to a sequence of increasingly prolonged wiggles in the amplitude. This behavior reflects the dispersive nature of the massive field and the evolving phase difference between overlapping wave components.
@@ 2 Schwarzschild black hole with the Gaussian bump in the effective potential
From the plot 1 one can see that, when the bump is located far from the black hole, in agreement with (2.12) echoes appear at very late times, where we observe the modification of the power-law tail envelope. When the bump is closer to the event horizon, as shown in Fig. 2, both the intermediate tails and the final stage of the ringing phase are significantly modified. However, the asymptotic late-time tails, being governed by the asymptotic behavior of the potential, change insignificantly. Therefore, in order to observe the distinct echoes for the massive field, we need a significant modification of the effective potential outside the radiation zone. Such a modification can appear at a distance, due to another compact object with a comparable mass, e.g., a black hole or a neutron star, or, alternatively, due to the effects of new physics near the horizon or instead of the horizon. The latter effect for a Schwarzschild wormhole we shall consider in Sec. 4.
For sufficiently small bumps near the event horizon, the time-domain profile during the ringdown phase does not differ significantly from that of the spacetime without a bump. This result aligns with [62, 63], where it was demonstrated that the fundamental mode typically changes only slightly when the potential is slightly deformed near the horizon, with stronger deviations seen in the first few overtones. These overtones decay quickly, and for massive fields, their contribution to the time-domain profile is suppressed because (a) the fundamental mode decays more slowly and (b) the tails of massive fields begin to dominate earlier, leaving a relatively short period for the ringing phase.
Figure 3: Semi-logarithmic plot of the time-domain profiles for a massive scalar field perturbations around the Schwarzschild black hole with a bump: $r_{0}=1$, $\ell=1$, $r_{m}=8$, $\kappa=1$, $\mu=1$ (blue) and $\mu=0$ (red), $A=1/100$ (top panel) and $A=1/12$ (bottom panel). The observer is situated at $r_{p}=3$, and the center of the Gaussian wave-package is between the observer and the event horizon, $p=16$. The black slopes show exponential decay of the echoes’ amplitudes: $\propto\exp(-0.02t)$ for $A=1/100$ (top) and $\propto\exp(-0.01t)$ for $A=1/12$ (bottom). The orange slope in the top panel shows the final signal after relaxation, governed by the quasinormal modes of the composite potential. The higher the bump, the later the time at which full relaxation occurs.
When comparing the echoes for massive and massless fields (see Fig. 3), we observe that the massless field, being long ranged, scatters the energy faster and, therefore the amplitude of echoes are many orders smaller. Even for an unrealistically, for astrophysical environment, large bump ($A=1/12$) we see that, although the amplitude of the first echo is of the same order of magnitude as the intermediate time tail, each next echo’s amplitude becomes by orders smaller than the echo for the massive field. The amplitudes of the sequence of echoes decrease exponentially. Eventually, after the signal has relaxed, we observe an exponential decay governed by the quasinormal modes of the full potential characterized by two peaks.
@@
Figure 4: Logarithmic plot of the time-domain profiles for a massive scalar field perturbations around the Schwarzschild black hole with a bump: $r_{0}=1$, $\ell=1$, $r_{m}=8$, $\kappa=1$, $\mu=2$, $A=1/100$ (left panel) and $A=1/12$ (right panel). The black slopes show power-law decay of the intermediate tails ($\propto t^{-2}$) and the orange slopes show power-law decay of the amplitudes of the echoes at late times ($\propto t^{-5/6}$).
In contrast, a massive field decays according to a power-law tail. At intermediate times, the power-law tail interacts with the echoes generated by the bump, resulting in oscillations in the signal’s amplitude. The period of these oscillations increases over time as the signal relaxes, eventually giving way to the universal asymptotic power-law tail decay for massive fields (see Fig. 4), given by Eq. (2.11).
Thus, perturbations of massive fields decay according to a power law at late times. As a result, the late-time amplitude of a massive field remains significantly higher (by several orders of magnitude) compared to that of a massless field. Even in the case of an unrealistically large bump, which enhances the amplitude of echoes, the exponential decay of massless fields continues into later times and ultimately results in a smaller amplitude compared to the late-time behavior of the massive field. This reflects the persistent power-law decay of the latter.
3 Gravitational perturbations of squashed black holes in the Kaluza-Klein theory
Squashed Kaluza-Klein (KK) black holes [64] exhibit significant differences from standard Schwarzschild black holes, which are characterized by asymptotic flatness, and from black strings, even at energy levels where KK modes have not yet been activated. This makes squashed KK black holes a valuable model for exploring the dynamics of higher-dimensional spacetimes. Unlike their more conventional counterparts, these squashed black holes offer unique insights into higher dimensions, providing a glimpse into physics beyond the familiar four-dimensional framework.
One particularly noteworthy aspect of squashed KK black holes is their apparent stability [24]. This distinguishes them from other higher-dimensional solutions, such as black strings, which are prone to the Gregory-Laflamme instability [65]—a phenomenon that can cause certain configurations to become dynamically unstable and fragment. The inherent stability of squashed KK black holes allows us to bypass this instability, making them more viable candidates.
The metric of the uncharged non-rotating squashed Kaluza-Klein black hole is [24]
$ds^{2}$ $=$ $-F(\rho)dt^{2}+\frac{G(\rho)^{2}}{F(\rho)}d\rho^{2}+\rho^{2}G(\rho)^{2}(\sigma_{1}^{2}+\sigma_{2}^{2})+\frac{\rho_{\infty}^{2}}{4G(\rho)^{2}}(\sigma_{3})^{2},$ (3.1)
$\sigma_{1}$ $=$ $-\sin{\psi}d\theta+\cos{\psi}\sin{\theta}d\phi\ ,$
$\sigma_{2}$ $=$ $\cos{\psi}d\theta+\sin{\psi}\sin{\theta}d\phi\ ,$ (3.2)
$\sigma_{3}$ $=$ $d\psi+\cos{\theta}d\phi\ ,$
and
$\rho_{\infty}^{2}=4\rho_{0}(\rho_{+}+\rho_{0})\ ,$
As shown in [66], the linearized perturbation equations of the gravitational field for zero modes with respect to the $SU(2)$ group are also reduced to the wave-like form with the following effective potentials,
$+64\rho^{3}(17\rho_{+}+2\rho)\rho_{0}+48\rho^{2}(32\rho_{+}+11\rho)\rho_{0}^{2}+60\rho(13\rho_{+}+12\rho)\rho_{0}^{3}+9(9\rho_{+}+35\rho)\rho_{0}^{4}\Big]\ ,$
$V_{1}$ $=$ $\left(1-\frac{\rho_{+}}{\rho}\right)\bigg[\frac{1}{\rho_{+}\rho_{0}}+\frac{7(-\rho_{+}+\rho)}{16(\rho+\rho_{0})^{3}}+\frac{17\rho_{+}-19\rho}{8\rho(\rho+\rho_{0})^{2}}+\frac{9(-3\rho_{+}+7\rho)}{16\rho^{2}(\rho+\rho_{0})}$
$+\frac{\rho_{+}-\rho}{\rho_{+}^{2}\rho+\rho_{+}\rho\rho_{0}}-\frac{8(\rho_{+}-\rho)^{2}\rho}{(\rho_{+}-2\rho)(\rho_{+}\rho_{0}-\rho(\rho+2\rho_{0}))^{2}}-\frac{2(\rho_{+}+2\rho)}{(\rho_{+}-2\rho)(\rho_{+}\rho_{0}-\rho(\rho+2\rho_{0}))}\bigg]\ ,$
$V_{2}$ $=$ $\frac{-(\rho_{+}-\rho)}{16\rho^{3}\rho_{0}(\rho_{+}+\rho_{0})(\rho+\rho_{0})^{3}}\Big[64\rho^{5}+256\rho^{4}\rho_{0}-32\rho^{3}\rho_{0}(\rho_{+}-11\rho_{0})$
$+9\rho_{+}\rho_{0}^{3}(\rho_{+}+\rho_{0})+8\rho^{2}\rho_{0}(2\rho_{+}^{2}-5\rho_{+}\rho_{0}+25\rho_{0}^{2})+\rho\rho_{0}^{2}(20\rho_{+}^{2}-9\rho_{+}\rho_{0}+35\rho_{0}^{2})\Big]\ .$